# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ H.3 Upper bounds for scheduled and aggregated updates (general $m$)
$\mathrm{UB}_{\text{agg}}(x;\eta):=F(x)\;-\;\eta\,\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\frac{L\eta^{2}}{2}\,\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}$ (H.11)
The integrals in Equation H.10 are over ordered pairs $p<q$ along the specific sequence $x_{0}\to x_{1}\to\cdots\to x_{m}$; the bound therefore depends on the within-refresh order. Randomizing the order yields an expected version.
In Sections H.4H.6 we express the scheduled bound in terms of $\{G_{r}^{0}\}$ under the per-group lipschitzness assumption. The associated drift terms are aggregated into $R_{m}(x;\eta)$.
H.4 Scheduled and aggregated gap at a common linearization
Define the shorthand
$I_{pq}(x;\eta)\ :=\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle\,dt$ (H.12)
Theorem 8 (Upper-bound gap under per-group lipschitzness).
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\;\leq\;\eta^{2}\!\sum_{1\leq p<q\leq m}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;+\;I_{pq}(x;\eta)\Big)\;+\;R_{m}(x;\eta).$ (H.13)
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz
$I_{pq}(x;\eta)\ \leq\ L\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|$ (H.14)
which gives the envelope
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ L\eta^{2}\!\sum_{p<q}\!\big(\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|-\langle G_{p}^{0},G_{q}^{0}\rangle\big)\;+\;R_{m}(x;\eta)\ \ \geq\ 0$ (H.15)
Interpretation
Proposition 9 (Drift penalty bound under per-group lipschitzness).
Assume each group map $G_{r}$ is $L_{r}$-lipschitz. Then for $r\geq 2$,
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\eta\sum_{p<r}\|G_{p}^{0}\|\ :=\ \|G_{r}^{0}\|+L_{r}\eta\,S_{r-1},$
and the scheduled start substitution error satisfies
$R_{m}(x;\eta)\ \leq$ $\ \eta^{2}\Big(\sum_{p=1}^{m}\|G_{p}^{0}\|\Big)\,\sum_{r=2}^{m}L_{r}\,S_{r-1}$
H.5 Sufficient conditions for a tighter scheduled bound
Assumption 7 (Hessian-weighted negative cross-terms).
There exist nonnegative margins $\{\Gamma_{pq}\}_{p<q}$ such that
$I_{pq}(x;\eta)\ =\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle dt\ \leq\ -\,\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\quad\text{for all }p<q$ (H.16)
Assuming per-group lipschitzness and H.16, for any $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ \eta^{2}\!\sum_{p<q}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;-\;\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\Big)\ +\ R_{m}(x;\eta)$ (H.17)
In particular, if
$\sum_{p<q}\!\Big(\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|+L\,\langle G_{p}^{0},G_{q}^{0}\rangle\Big)\ >\ \frac{R_{m}(x;\eta)}{\eta^{2}}$ (H.18)
then $\mathrm{UB}_{\text{sch}}(x;\eta)\ <\ \mathrm{UB}_{\text{agg}}(x;\eta)$
H.6 PL or strong convexity: standard rate and upper-bound gains for scheduling
$\tfrac{1}{2}|\nabla F(x)|^{2}\ \geq\ \mu\,(F(x)-F^{\star}),\qquad\forall x$ (H.19)
$F(x^{\text{agg}})\ \leq\ F(x)\;-\;\eta\Big(1-\tfrac{L\eta}{2}\Big)\,|\nabla F(x)|^{2}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)$ (H.20)
$\Delta{\text{UB}}(x;\eta)\ :=\ \mathrm{UB}_{\text{agg}}(x;\eta)-\mathrm{UB}_{\text{sch}}(x;\eta)\ \geq\ 0$ (H.21)
whenever H.18 holds. Since $F(x^{\text{sch}})\leq\mathrm{UB}_{\text{sch}}(x;\eta)$ and $\mathrm{UB}_{\text{agg}}(x;\eta)$ upper-bounds the one-shot decrease term in H.20, we obtain the bound-level contraction
$F(x^{\text{sch}})-F^{\star}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)\;-\;\Delta_{\text{UB}}(x;\eta).$ (H.22)
Consequently, under per-group lipschitzness and H.18, the scheduled refresh satisfies the standard gradient-descent contraction and, in addition, achieves an extra nonnegative decrement $\Delta_{\mathrm{UB}}(x;\eta)$ in the upper bound.
H.7 Why the assumptions are mild
H.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality [44, 91].
H.7.2 Per-group lipschitzness of $G_{r}$
H.7.3 Negative Hessian-weighted cross-terms
H.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use ([92, 93, 94]).
H.8 Concluding remarks
[…]

### v2

@@ Appendix F Convergence rate with $\tau$ -dependent constant
$\tau_{\mathrm{eff}}(S):=\frac{\sum_{i\neq j\in S}\bigl(-\langle g_{i},g_{j}\rangle\bigr)_{+}}{\sum_{k\in S}\|g_{k}\|^{2}}\in[0,\infty),\qquad(x)_{+}=\max\{x,0\}.$ (93)
Then for every step $t$,
$\Bigl\|\textstyle\sum_{k\in S_{t}}g_{k,t}\Bigr\|^{2}\;\geq\;\Bigl(1-\tau_{\mathrm{eff}}(S_{t})\Bigr)\sum_{k\in S_{t}}\|g_{k,t}\|^{2}.$ (94)
Consequently,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\underbrace{\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\frac{1}{1-\tau_{\mathrm{eff}}(S_{t})}\Bigr]}_{=:~\Gamma_{T}}\cdot\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\|g_{t}\|^{2}\bigr].$ (95)
Using $\mathbb{E}\|g_{t}\|^{2}=\mathbb{E}\|\nabla F(\theta_{t})\|^{2}+\mathbb{E}\|g_{t}-\nabla F(\theta_{t})\|^{2}\leq\mathbb{E}\|\nabla F(\theta_{t})\|^{2}+\sigma^{2}$ and the average version of equation 91,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\|\nabla F(\theta_{t})\|^{2}\bigr]\;\leq\;\frac{2(F_{0}-F^{\star})}{\eta T}\;+\;L\eta\sigma^{2},$ (96)
we obtain the $\tau$-dependent, data-driven control
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\Gamma_{T}\Biggl(\frac{2(F_{0}-F^{\star})}{\eta T}+L\eta\sigma^{2}+\sigma^{2}\Biggr).$ (97)
If, in addition, each $S_{t}$ is pairwise $\tau$-compatible with $|S_{t}|=s_{t}$ and $\tau\,(s_{t}-1)\leq\rho<1$ uniformly in $t$, then $\tau_{\mathrm{eff}}(S_{t})\leq\tau(s_{t}-1)\leq\rho$ and hence $\Gamma_{T}\leq\tfrac{1}{1-\rho}.$ With $\eta=\tfrac{c}{\sqrt{T}}$, equation 97 becomes
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\frac{1}{1-\rho}\left(\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}\;+\;\sigma^{2}\right).$ (98)
F.1 Discussion and intuition
Equation equation 90 is the classical $O(1/\sqrt{T})$ rate for non-convex SGD with unbiased and bounded variance gradients and constant-over-time step size $\eta=c/\sqrt{T}$. Under these conditions, the convergence rate in terms of the full gradient norm $\|\nabla F(\theta_{t})\|^{2}$ does not depend on $\tau$. However, the scheduler’s $\tau$ structure does control the per step energy of the scheduled gradient through equation 95–equation 98. Less cross-task conflict (smaller $\Gamma_{T}$) results in a tighter bound on $\tfrac{1}{T}\sum_{t}\sum_{k\in S_{t}}\|g_{k,t}\|^{2}$, which is the quantity governed by the descent preservation inequalities used throughout the analysis.
Appendix G Bounded Staleness via Greedy Graph Coloring
Proposition 8 (Staleness Bound).
$m\;\leq\;\Delta+1.$ (99)
$s_{\max}\;=\;m-1\;\leq\;\Delta$ (100)
iterations. In particular, the schedule enforces a bounded inter-update delay of at most $\Delta$ iterations per task, consistent with the bounded-delay assumption of Recht et al. (Niu et al., 2011).
Proof.
We proceed in two parts.
Part A: Color count bound.
A greedy algorithm scans vertices in some order and assigns to each vertex the smallest available color not used by its already colored neighbors. When the $i$-th vertex $v$ is reached, at most $\deg(v)\leq\Delta$ of its neighbors are already colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ (Lovász, 2006).
Part B: Staleness of cyclic execution.
Fix any task $T\in\mathcal{T}$ and let it belong to color $C_{j}$ for some $1\leq j\leq m$. Under cyclic scheduling, $C_{j}$ is executed at steps $t=j,\;j+m,\;j+2m,\dots$. The number of intervening steps between two consecutive executions of $C_{j}$ is exactly m1m-1. Hence task $T$ never waits more than $s_{\max}=m-1$ iterations for an update. Combining with Equation 8 yields $s_{\max}\leq\Delta$. ∎
G.1 Interpretation
The bound (Equation 100) guarantees that the shared parameters used by any task are refreshed at least once every $\Delta$ iterations in the worst case (e.g., when the conflict graph is a clique of size $\Delta+1$). This aligns with the bounded-delay assumption common in analyses of asynchronous SGD and lock-free training, so convergence proofs built under that assumption apply to our cyclic schedule with delay parameter at most $\Delta$ when iterations are used as the unit of delay (Niu et al., 2011; Lian et al., 2015). In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
Appendix H Greedy Graph-Coloring Uses at Most $\Delta\!+\!1$ Colors
H.1 Proof of Proposition 9
Proposition 9 (Coloring Period Bound).
Let $G=(V,E)$ be a finite, simple, undirected graph with maximum degree $\Delta\,\coloneqq\!\max_{v\in V}\deg(v)$. The greedy (first-fit) coloring algorithm (e.g., Welsh–Powell order) [^1] Order the vertices in non–increasing degree and assign to each the smallest positive integer (color) not used by its previously colored neighbors. produces a proper vertex coloring with no more than
$\chi_{\text{greedy}}(G)\;\leq\;\Delta+1$ (101)
distinct colors. Consequently, when the scheduler activates the color classes in a cyclic order, the cycle length is bounded by $\Delta+1$. This is a quantity depending only on the structure of the conflict graph.
Proof.
Let the vertices be processed in the chosen order $v_{1},v_{2},\dots,v_{|V|}\,$ (e.g., Welsh-Powell). Assume inductively that after coloring the first k1k-1 vertices the algorithm has used at most $\Delta+1$ colors. Consider vertex $v_{k}$. Since $\deg(v_{k})\leq\Delta$, at most $\Delta$ neighbors of $v_{k}$ can appear before $v_{k}$ in the ordering. Hence, at the moment of coloring $v_{k}$, at most $\Delta$ colors are forbidden (one for each previously colored neighbor). Among the palette $\{1,2,\dots,\Delta+1\}$ there is therefore at least one color still available. Assigning the smallest such color to $v_{k}$ maintains a proper coloring and never introduces a new color beyond $\Delta+1$.
Proceeding vertex-by-vertex, no step ever requires more than $\Delta+1$ colors, establishing equation 101. ∎
H.2 Implications for the scheduler
A coloring with at most $\Delta+1$ classes means the scheduler’s cycle period (the number of batches needed before every task reappears) is bounded by a graph invariant independent of the number of tasks. Even if thousands of tasks exist, as long as each one conflicts with at most $\Delta$ others, the memory footprint (one shared backbone plus $\Delta+1$ sets of head activations) and the maximum waiting time between successive updates for any task (bounded by $\Delta$, see Proposition 8) remain predictable and small. This guarantee is essential for scaling the scheduler to large, heterogeneous tasks.
Appendix I Baseline Non-Convex SGD Convergence Rate
I.1 Proof of Theorem 8
Theorem 8 (Classical $O(1/\sqrt{T})$ bound).
Let $F:\mathbb{R}^{d}\!\to\!\mathbb{R}$ be an $L$-smooth, possibly non-convex objective and suppose the stochastic gradient $g_{t}$ computed at iteration $t$ satisfies
$\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t}),\qquad\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\;\leq\;\sigma^{2}.$ (102)
$\min_{0\leq t<T}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2\bigl(F_{0}-F^{\ast}\bigr)}{c\sqrt{T}}+\frac{cL\sigma^{2}}{\sqrt{T}},$ (103)
where $F^{\ast}=\inf_{\theta}F(\theta)$.
Proof.
The proof is a streamlined restatement of ((Ghadimi & Lan, 2013; Nemirovski et al., 2009)). By $L$-smoothness,
$F(\theta_{t+1})\;\leq\;F(\theta_{t})+\langle\nabla F(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\lVert\theta_{t+1}-\theta_{t}\rVert^{2}.$ (104)
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\leq\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\eta\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L}{2}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr].$ (105)
Decompose the squared stochastic gradient:
$\mathbb{E}[\lVert g_{t}\rVert^{2}]=\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\sigma^{2}$ (106)
Thus, and using $\eta\leq 1/L$ so that $\eta-\tfrac{L\eta^{2}}{2}\geq\tfrac{\eta}{2}$,
$\mathbb{E}[F(\theta_{t+1})]\leq\mathbb{E}[F(\theta_{t})]-\frac{\eta}{2}\,\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (107)
Summing from $t=0$ to T1T-1 and telescoping gives
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]\leq F_{0}-F^{\ast}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (108)
Dividing by $\eta T$ and inserting $\eta=c/\sqrt{T}$ yields equation 103. ∎
I.2 Connection to the scheduler
At $\tau=0$, pairs with negative inner product are incompatible, so the conflict graph on tasks can be colored into $m$ classes $\{C_{1},\dots,C_{m}\}$, and a simple policy activates one color class per step. Under a deterministic (cyclic) activation order, the update $g_{t}=\sum_{k\in S_{t}}g_{k,t}$ generally satisfies
$\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\sum_{k\in S_{t}}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})\;\neq\;\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t}),$ (109)
so it is biased for the full gradient.
I.2.1 Consistency with the unbiased SGD assumption
The analysis in Theorem 8 assumes an unbiased stochastic gradient, $\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t})$. This assumption is met under either of the following implementations.
(i) Randomized class sampling with scaling. Draw $J_{t}\sim\mathrm{Unif}\{1,\ldots,m\}$ independently each step and set
$\tilde{g}_{t}\;=\;m\!\!\sum_{k\in C_{J_{t}}}\!g_{k,t}.$ (110)
Then $\mathbb{E}[\tilde{g}_{t}\mid\theta_{t}]=\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})=\nabla F(\theta_{t}),$ so Theorem 8 applies (with the variance bound adjusted for the scaled estimator). Equivalently, one may keep $g_{t}=\sum_{k\in C_{J_{t}}}g_{k,t}$ and use an effective step size $m\eta$.
(ii) Deterministic cyclic schedule. If the classes are visited in a fixed periodic order, then generally $\mathbb{E}[g_{t}\mid\theta_{t}]\neq\nabla F(\theta_{t})$ at the per-step level. Nonetheless, standard analyses of nonconvex smooth cyclic block updates yield an $O(1/\sqrt{T})$ decay of the average gradient norm under usual step-size conditions, with constants depending on the number of blocks.
Either implementation delivers an $O(1/\sqrt{T})$ convergence guarantee.
Appendix J Experimental Setup for Datasets
Table 3: Information on the datasets utilized in experimentation. (*Some samples were removed during preprocessing)
Dataset Main Tasks (+) Aux. Tasks (-) Aux Tasks Modalities Samples
AV-MNIST Digit Classification Digit Parity Audio, Image 56.0k
$M$-IMDb Genre Classification Release Decade Title-Iniial Classification Image, Text 25.9k
We evaluate the proposed scheduler alongside numerous baselines and state-of-the-art models across multiple datasets to reliably assess its general performance relative to other approaches. In total, it is evaluated across 6 datasets.
Across all datasets, we incorporate positive and/or negative auxiliary tasks into training. Positive auxiliary tasks share structure or predictive signals with the main tasks (e.g., common features or correlated outputs) and so can improve the learned representations by providing relevant supervision. In contrast, negative auxiliary tasks are uncorrelated or directly conflicting with the main objectives, inducing gradient interference that can slow or degrade primary performance. Including both creates controlled variation in task alignment, letting us test whether SON-GOKU (1) groups compatible tasks, (2) separates conflicting tasks, and (3) maintains main-task performance under interference created by auxiliary tasks.
J.1 NYUv2
The NYU Depth Dataset v2 (NYUv2) (Silberman et al., 2012) consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 (He et al., 2015) backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
J.2 CIFAR-10
The CIFAR-10 (Krizhevsky et al., 2009) dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite (Hendrycks & Dietterich, 2019), and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 (He et al., 2015) backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
J.3 AV-MNIST
The AV‑MNIST benchmark (Vielzeuf et al., 2018) pairs MNIST images (Lecun et al., 1998) with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS (Leonard & Doddington, 1993). It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion (Boulahia et al., 2021; Guarrasi et al., 2025), as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets (Tacchetti et al., 2018; Mohammadi et al., 2020).
J.4 MM-IMDb
The MM‑IMDb dataset (Arevalo et al., 2017) contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 (Simonyan & Zisserman, 2014) and Google word2vec (Mikolov et al., 2013) model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
J.5 STOCKS
The STOCKS datasets we use, introduced in (Liang et al., 2021), contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category (MSCI Inc. & S&P Dow Jones Indices, 2024), and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM (Cui et al., 2018).
Appendix K Models Used for Comparison
K.1 Baseline models
  1. 1.
  2. 2.
    Gradnorm (Chen et al., 2018). Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
  3. 3.
    MGDA (Sener & Koltun, 2018). Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions (Lockwood, 2008; Pareto, 2014) through gradient descent in the convex hull of gradients (Fliege & Svaiter, 2000; Miettinen, 1999).
K.2 State-of-the-art models
  1. 1.
    PCGrad (Yu et al., 2020). Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
  2. 2.
    CAGrad (Liu et al., 2021). Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
  3. 3.
    Adatask (Yang et al., 2023). Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
  4. 4.
    FAMO (Liu et al., 2023). Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
  5. 5.
    Fair Resource Allocation in MTL (FairGrad) (Ban & Ji, 2024). Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
  6. 6.
    Nash-MTL (Navon et al., 2022). Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
Appendix L Expanded Wall-Clock Time Study
Table 4: We present wall-clock time (seconds $\pm$ standard deviation) across all $K$ and scheduler refresh rates $R\in\{4,32,256\}$. We split results into sub-tables by $R$ for readability. Non-scheduler methods do not depend on $R$, so they are shown in the $R=4$ sub-table and omitted in the R=32,256R{=}32,256 subtables to avoid redundancy.
(a) R=4R{=}4 (all methods)
Method K=3 K=6 K=16 K=40
Uniform 0.2656 $\pm$ 0.1201 0.3240 $\pm$ 0.0629 0.3798 $\pm$ 0.1050 0.4054 $\pm$ 0.1190
GradNorm 5.4714 $\pm$ 0.7137 5.1201 $\pm$ 0.6112 4.9042 $\pm$ 0.5869 4.7372 $\pm$ 0.9286
MGDA 97.1081 $\pm$ 5.4645 121.4371 $\pm$ 9.0923 132.4913 $\pm$ 3.1752 134.0878 $\pm$ 2.2621
PCGrad 3.6212 $\pm$ 0.3517 23.1266 $\pm$ 0.8773 176.7566 $\pm$ 2.8171 1127.1337 $\pm$ 34.2603
CAGrad 102.8651 $\pm$ 18.3422 136.1034 $\pm$ 2.4218 134.3585 $\pm$ 4.0791 132.7034 $\pm$ 1.2412
AdaTask 2.1816 $\pm$ 0.0934 2.1032 $\pm$ 0.1012 2.2853 $\pm$ 0.0718 2.2278 $\pm$ 0.1370
FAMO 2.0725 $\pm$ 0.2073 1.9980 $\pm$ 0.1998 2.1710 $\pm$ 0.2171 2.1164 $\pm$ 0.2116
FairGrad 3.8020 $\pm$ 0.5703 15.2079 $\pm$ 2.2812 108.1450 $\pm$ 16.2218 675.9065 $\pm$ 101.3860
Nash-MTL 5.7030 $\pm$ 1.1406 22.8118 $\pm$ 4.5624 162.2176 $\pm$ 32.4435 1013.8598 $\pm$ 202.7720
SON-GOKU 2.0904 $\pm$ 0.3506 3.6770 $\pm$ 0.4974 6.3225 $\pm$ 0.7895 14.3280 $\pm$ 1.4073
SON-GOKU ++ AdaTask 4.1011 $\pm$ 0.4174 5.2126 $\pm$ 0.6066 7.6798 $\pm$ 0.7107 14.7528 $\pm$ 1.8671
SON-GOKU ++ GradNorm 7.3223 $\pm$ 0.4994 8.5898 $\pm$ 0.8203 12.1065 $\pm$ 2.5850 16.8329 $\pm$ 1.9803
SON-GOKU ++ PCGrad 2.3489 $\pm$ 0.3258 3.5925 $\pm$ 0.4100 6.1549 $\pm$ 0.8461 12.5729 $\pm$ 1.2657
(b) R=32R{=}32 (scheduler-based approaches)
Method K=3 K=6 K=16 K=40
SON-GOKU 1.9896 $\pm$ 0.3651 3.3202 $\pm$ 0.5745 6.0897 $\pm$ 0.9425 12.1432 $\pm$ 1.2044
SON-GOKU ++ AdaTask 3.7718 $\pm$ 0.9654 5.0511 $\pm$ 0.6531 7.5903 $\pm$ 1.1920 14.5182 $\pm$ 2.0660
SON-GOKU ++ GradNorm 7.0202 $\pm$ 1.0711 8.1661 $\pm$ 0.9355 10.7227 $\pm$ 2.2088 16.5760 $\pm$ 1.8418
SON-GOKU ++ PCGrad 1.9834 $\pm$ 0.3586 3.4971 $\pm$ 0.3840 6.1395 $\pm$ 0.9425 10.9097 $\pm$ 1.5263
(c) R=256R{=}256 (scheduler-based approaches)
Method K=3 K=6 K=16 K=40
SON-GOKU 1.7593 $\pm$ 0.2280 3.0024 $\pm$ 0.3942 4.8411 $\pm$ 0.7302 11.4162 $\pm$ 1.6076
SON-GOKU ++ AdaTask 3.7224 $\pm$ 0.2696 4.4548 $\pm$ 0.5837 7.5276 $\pm$ 0.6230 13.0608 $\pm$ 3.2925
SON-GOKU ++ GradNorm 6.0221 $\pm$ 1.0418 7.8659 $\pm$ 0.7917 9.5029 $\pm$ 1.2168 15.6860 $\pm$ 2.3680
SON-GOKU ++ PCGrad 1.6776 $\pm$ 0.4104 3.0189 $\pm$ 0.7854 5.9893 $\pm$ 1.3797 7.1915 $\pm$ 0.2021
We provide more results from our wall-clock time study. The expanded table includes results from testing refresh rates $R\in\{4,32,256\}$ for scheduler-based methods.
L.1 Experimental Setup for Wall-Clock Time Study
We benchmark wall-clock time with a controlled synthetic workload to remove the effects of data loading and I/O. For each configuration (number of tasks $K$ and scheduler refresh rates $R$), we pre-generate a fixed sequence of per-task gradient vectors and loss values directly on the target device, and then feed the same exact tensors, in the same exact order, to every method. We set the gradient dimensionality to 1024. Timing uses a high-resolution clock with a device synchronize before starting and after finishing to capture only on-device compute. We also accumulate the norm of the combined gradient into a scalar accumulator (also known as a scalar sink) so the backend must realize the computation, avoiding lazy evaluation. Each MTL approach is run for 900 steps and repeated 10 times.
Appendix M Extended Related Work
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation (Yang et al., 2023). Early adaptive-weighting approaches sought to balance losses automatically (Vandenhende et al., 2022; Fan et al., 2023), while more recent work modifies gradients directly (Yu et al., 2020). Task scheduling and grouping methods, though far less popular than adaptive weighting techniques (Torbarina et al., 2023), have contributed to the field by controlling the timing of updates.
M.1 Tuned Loss Weighting
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others (Kurin et al., 2022; Zhao et al., 2024; Mueller et al., 2022), especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning (Argyriou et al., 2007; Ando & Zhang, 2005; Evgeniou et al., 2005; Kang et al., 2011; Liang & Zhang, 2020; Lin et al., 2022; Yu et al., 2021), but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
M.2 Adaptive Loss Weighting
(Kendall et al., 2018) introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) (Bishop, 2006) noise to scale losses automatically and improve depth and semantics on NYUv2 (Silberman et al., 2012).
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds (Chen et al., 2018). It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is scaled. This eliminates the need for expensive grid searches over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks (Liu et al., 2019).
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks (Yang et al., 2023).
M.3 Gradient-Level Conflict Mitigation
[…]